two circles of radius 10 cm are drawn so that their centers are 12 cm apart. The two points of intersection determine a common chord. Find the length of this chord.
thx
: two circles of radius 10 cm are drawn so that their centers are 12 cm apart. The two points of intersection determine a common chord. Find the length of this chord.
thx

Answer by Alan3354(1218)

two circles of radius 10 cm are drawn so that their centers are 12 cm apart. The two points of intersection determine a common chord. Find the length of this chord.
-------------
If you draw them, you'll see that there is a triangle formed by the 2 centers and one of the points of intersection of the circle. The chord is perpendicular to the line connecting the 2 circles, so it makes a right triangle with the radius and 1/2 of the line between the centers.
The radius is the hypotenuse = 10.
1/2 of the line connecting the centers = 6.
So 1/2 the chord = sqrt(10^2 - 6^2)
C/2= sqrt(64)
C/2 = 8
The chord = 16.